b).For the past records, about 40% of a firms orders are for export . Their record for export is 48% in one particular financial quarter. If they expect to sastify about 80 orders in the next financial quarter, what is the probability that they will break their previous export record?.
1
Expert's answer
2020-02-19T09:53:41-0500

For large values of n, the distributions of the count X and the sample proportion "\hat{p}" are approximately normal. In practice, the approximation is adequate provided that both "np\geq5" and "n(1-p)\geq10" , since there is then enough symmetry in the underlying binomial distribution.

Given the proportion "p" is expected to be 0.48 and from past records "\hat{p}" was found to be 0.4 and "n=10." Then

"np=10(0.48)=4.8<5,"

"n(1-p)=10(1-0.48)=5.2<10"

Normal distribution cannot be used.

Binomial distribution is appropriate for this event. Then "10(0.4)=4"

Probability of breaking the record

"P(X>4)=1-P(X=0)-P(X=1)-" "P(X=2)-P(X=3)-P(X=4)="

"=1-\dbinom{10}{0}(0.48)^0(1-0.48)^{10-0}-\dbinom{10}{1}(0.48)^1(1-0.48)^{10-1}-"

"-\dbinom{10}{2}(0.48)^2(1-0.48)^{10-2}-\dbinom{10}{3}(0.48)^3(1-0.48)^{10-3}-"

"-\dbinom{10}{4}(0.48)^4(1-0.48)^{10-4}\approx"

"\approx1-0.0015-0.0133-0.0554-0.1364-0.2204\approx"

"\approx0.573"